Levine, Inc., has a total debt ratio of .43. What is its debt-equity ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Debt-equity ratio What is its equity multiplier? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Equity multiplier
Answer #1

Answer:

Debt ratio = Total debt/Total asset

0.43 = Total debt TD)/ (Total debt + Total equity (TE))

0.43TD + 0.43TE = TD

0.43TE = 0.57TD

TD/TE = 0.43/0.57

            = 0.75

Debt equity ratio = 0.75

Equity multiplier = 1+ TD/TE

                                = 1+0.75

                                = 1.75